Sol Summary – November 12th

MDRS Daily Summary Report for sol 0.9

Summary Title: It’s sunday.

Mission Status: Today looks like a really Sunday. The crew is more relaxed but also focused with their duties. We had a good meal today prepared by Camila. Some of us shared some activities at this time, preparing the lunch. We are a team!

Sol Activity Summary: Keeping a High State of Consciousness in MDRS.

Look Ahead Plan: Tomorrow, We are planning an EVA walking to Lith Cannyon.

Anomalies in work:  Nope.

Weather: 9o C, 49o F

Crew Physical Status: Very well.

EVA: We did an interesting EVA without ATV. We visited an interesting spot walking to the north direction from the MDRS. Distance: 1,2 km. Was collected soil and rock samples. Was possible to do some beautiful images from the area.

Reports to be file: We sent in CapComm.

Support Requested: Nope.

EVA Report – November 12th

EVA #5 Report Sol 9

We began our EVA at 10:40 AM today. We started a bit late because we needed more time to prepare Carmen’s flight suit and test our communication equipment before this EVA. Since we didn’t need to use the rovers today, Julio and Carmen tried the Exploration suits. They didn’t seem particularly more comfortable as compared to the normal Suits. We walked towards the North Ridge area, until I could find a slope on the western slope that was shallow enough for us to climb. Once we found a shallow enough slope, we took samples from different levels of the hill. We took four samples, three from light-dark purple layers of regolith and one from a rusty-colored layer. We took several photographs of the surrounding landscape as well as the sample sites for later reference.

No falls or accidents on this outing. Sample Acquisition was successful! We began recompression at 11:55 PM. Distance covered was about 1.2-1.5 km total. Sample collection and ground covered went much quicker than expected. It’s important to note how quickly the shoulders and upper back can fatigue, even though we were only out for over an hour.

Prepared by Brandon Ferguson

GreenHab Report – November 12th

Sol 9 – 12/11/2017

Green Hab Officer Atila Meszaros

Ambient: heating 50 F

Ambient with window/door open: door opened at 11:32; closed at 17:00

Shade cloth on/off: on

Average temperature:

Time 08:40 13:41 15:36 18:46
Temperature inside (°C) 23 27 21 22
Temperature outside (°C) 2 13 12 8
% Humidity 16 16 16 16

Changes to crops:  Crops keep growing! A lot.

Daily water usage for crops: 2 galons

Time(s) of watering for crops: 08:40 and 15:36

Changes to research plants: first day of germination for amaranthus control! Finally!

Daily watering and amount of water used: 1 ½ L approximately

Narrative:

Hi, Jennifer! Atila in here, yes, like the Hun. I hope you are having an amazing night

The crops from a former experiment that is in the Green Hab (GH) keep growing. There is a gigantic bean seedling! It looks amazing. However, I’m worried about the tomato seedlings, they are too many in the same pot. I hope Shannon give me some instructions regarding this issue.  I elevated the temperature of the GH to 50 F.

Species (common name) Number of bowls Number of seedlings
Tomatoes 25 43
Cucumber 7 3
Melon 8 8
Beans 17 14
Pepper 9 3

Tomorrow as soon as possible I’ll update this numbers. It will be probably easier if I make a Drive with the dates and number of emerging seeds. Let me know if you have like a Data Base where I can put this kind of data.

Respecting my project. Half of the seeds remain at the GH; the other half is in the growing chamber. For all the seeds of the experiment, I’m using almost a Liter.

The project wants to prove that Peruvian native crops can, indeed, germinate and grow on Mars soil. I’m working with quinoa, amaranthus, and mustard (as a control).

The analog soil I’m using is becoming clay, and I’m worry that the seeds are not going to be able to break the soil and emerge. So, I’m going to germinate the seeds on Petri dishes and observe growing behavior on those ones. I already started today,

I think that’s pretty much all if you need more information about the GH I will be glad to send it to you.

Thank you for being there, Jenn!

Have a beautiful night on Earth.

Support/supplies needed: Distilled water (not extremely necessary for now)

Commander Report – Nov 12th

4h – Awake.

4h50m – 7h10m – 140 minutes walking: 12 km or 7,5 miles. Was beautiful to see so many stars in the night. I really like deserts during the night. It is a gift woke up early and have an experience like this.

Today, when I woke up, was 0o or 32oC. Walking in 0 or -2oC is an expression of resilience. When you are focused, there is no pain. This is a characteristic that must be considered in the astronauts experience.

I’m very thankful about the opportunity to prepare myself for this mission and keep the exercises during the simulation in MDRS.

9h – Breakfast of tortillas and cheddar cheese with coffee and

10h10m – 12h – EVA – I considered that We walked from the EVA and returning doing a distance around 2,5 km.  Our crew, always When We are in the decomprehension  chamber, to start an Extravehicular Activities, part of the crew who stays in the Hab plays a music to present an idea of time. The music of decompression was Coming Home of Iron Maiden. Interesting have a music that talks about aerospace.

During the EVA, I used the exploration spacesuit prototype. This spacesuit presents a painful ergonomics. It is hard have movements to collect rocks and soil samples.

13h – I’m doing my commander report. I organized my room and helped in the lunch preparation.

14h – 15h – Lunch.

16h – 17h30m – I took a nap. I was tired about 16 km or 10 miles walking today.

17h30m – From Brazil, I received news about the transmission of the interview of TV Globo related to the research associating Mars, MDRS and Brazilian semiarid. Some images from the visit in MDRS:

http://g1.globo.com/fantastico/edicoes/2017/11/12.html?utm_source=email&utm_medium=share-bar-desktop&utm_campaign=share-bar&#!v/6283988

17h30m – During the afternoon, after the lunch, Atila, Brandon and Camila worked in their research. I took some photos about sent in Crew-photos e-mail.

19h – CapComm.

Operations Report – November 11th

Crew 182 Operations Report  11Nov2017

SOL: 8

Name of person filing report:  Carmen Atauconcha

Non-nominal systems:

WATER HEATER: It is not working.  In the morning, we had to take a shower with cold water. The water heater is not heating the water.  We opened the entire water pipe, but it still did not heat.

COMMUNICATION RADIOS: The constantly noise of one of the radios stops.
SECONDARY AIRLOCK: The door of the secondary airlock is gotten out-off-square.  I am saying that it is wonky since we arrived on Mars.
ATV’s: The blue ATV is not working. When we try to turn it on, it does not answer.

Generator—  It worked the whole night since 6pm of 10Nov2017 until 8:00 am of 11Nov2017.  Turn on at 5.15 pm 11Nov2017.  Operand.

Solar—

READING FROM 11 NOV 2017

 

TIME SOC READINGS
9:11 am 100%
10:20 am 100%
11:20 am 100%
12:20 am 100%
13:20 am 100%
14:20 pm 100%
15:50 pm 98%
16:20 pm 97%
16:45 pm 94%
17:04 pm 91%

 

Diesel – 55 %

Propane – 50 psi

Gasoline (5 Gallon containers for ATV) – 1/4 gallons. In addition the tank of gasoline inside the ATV’s is filled.

Water (trailer) –  0 gallons

Water (static) – 258  gallons

Trailer to Static Pump used –   no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  36.6 gallons

Toilet tank emptied:  yes

ATV’s Used: No

Oil Checked:  No

Atv # Fuel Used Gals:  0  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 0 hours

Deimos used:  No

Hours: 0

Beginning charge:  –

Ending charge:  –

Currently charging: Yes

HabCar used? No

Notes and Comments:
Today we had an advance in the topic about recycling water. The number of gallons of water that we used today is less than the gallons of water used yesterday.  And, I know that tomorrow is going to be better.
Summary of  Hab operations:

We receive a new digester for the toilet.  And we wait that it removes the bad smell of the toilet. We will keep you post.

The area of the tools in the Hab was a mess. Today we clean and organize this place.  Also, in the afternoon, we flushed the toilet with the new digester.   From another hand, we are going to receive the communication radios on Monday.
Finally, today we received some food.  This is great for us. Thanks¡

Summary of GreenHab operations:

We turned on the heater at 18:12 pm with a temperature of 47 °F.  At the same time, we are cleaning this area to make it beautiful.

Summary of ScienceDome operations:

We received a new shop vac. It has been installed in the laboratory. Also, there were cables going through the laboratory that we order to keep the security in the laboratory.

Summary of health and safety issues:

The HOS officer needs two fire alarms and lamps for the science dome and for the greenhouse. In the morning we fixed the stair of the secondary airlock.  Now, that stair is more stable.

Questions, concerns and requests to Mission Support:

We need you to repair the heater water, since it still does not work.

For the next week at the most on Monday, we are going to need water.

I have a doubt, which is the minimum limit for the propane tank and the diesel tank?

Where does the used water go?

When are we going to receive water?

GreenHab – November 11th

Sol 8 – 11/11/2017

Green Hab Officer Atila Meszaros

Ambient: heating 45 F , turned on at 18:12

Ambient with window/door open: door opened at 08:55; closed at 17:39

Shade cloth on/off: on

Average temperature:

Time 09:00 13:00 17:05 18:42
Temperature inside (°C) 26 34 22 13
Temperature outside (°C) 3 6 11 1
% Humidity 16 16 16 16

 

Changes to crops:  Second day of germination! Crops keep growing!

Daily water usage for crops: 2 galons

Time(s) of watering for crops: 09:00 and 13:00

Changes to research plants: none.

Daily watering and amount of water used: 1 L approximately

Narrative:

Hi, Gael! Atila in here, like the Hun. I hope you are having a terrific time.

Today was the day after the germination of the crops from a former experiment that is on the Green Hab (GH).  The seeds keep germinating and the seedling keep growing.  I turned on the heater (45 F).

Species (common name) Number of bowls Number of seedlings
Tomatoes 25 8
Cucumber 7 1
Melon 8 7
Beans 17 7
Pepper 9 3

Tomorrow as soon as possible I’ll update this numbers. It will be probably easier if I make a Drive with the dates and number of emerging seeds. Let me know if you have like a Data Base where I can put this kind of data.

Respecting my project. Half of the seeds remain at the GH; the other half is in the growing chamber. For all the seeds of the experiment I’m using almost a Liter.

The project wants to prove that Peruvian native crops can, indeed, germinate and grow on Mars soil. I’m working with quinoa, amaranthus and mustard (as a control).

The analog soil I’m using is becoming clay, and I’m worry that the seeds are not going to be able to break the soil and emerge. So, I’m going to try a new procedure: I’m going to germinate the seeds on Petri dishes and observe growing behavior on those ones.

I think that’s pretty much all, if you need more information about the GH I will be glad to send it to you.

Thank you for being there!

Have a beautiful night on Earth.

Support/supplies needed: none

Sol Summary – November 11th

MDRS Daily Summary Report for sol 0.8

Summary Title: Reflecting about sustainability.

Mission Status: Our day was concentrated in search to prepare our mission plan, reflects about each individual research plan, and evaluates what kind of meals to be done.

Sol Activity Summary: improving the MDRS.

Look Ahead Plan: Tomorrow, We are planning an EVA walking close to the station.

Anomalies in work:  Nope.

Weather: 10o C, 50o F

Crew Physical Status: Fine.

EVA: No EVAs today.

Reports to be file: We sent in CapComm.

Support Requested: Nope.

Commander Report – November 11th

Yesterday night We had a meeting to clarify some points of our mission and try to improve our results in the next week in MDRS.

6h30m – I woke up.

6h42 – Planning.

6h53m – 8h23m – Walking 90 minutes: 8,6 km or 5,3 miles.

9h10m – Breakfast.

10h20m – Recycling some plastic bottles, and trying to fix something.

12h30m – Lunch: eggs and Hash brown potatoes.

13h – We debated about the mission plan, considering about each research plan. I consider that the main topic of our mission is the sustainability.

15h45m – Camila and Brandon prepared a banana cake.

16h22m 17h02 – Walking 40 minutes: 5,8 km or 3,6 miles. At total I walked today 130 minutes.

Today didn’t happened EVA.

17h20m – I’m working in my research related with sustainability in MDRS:

Sustainability in MDRS:

lessons to Mars research habitats and setlements

Researcher: Prof. Julio Francisco Dantas de Rezende, PhD.

This research searches to answer the question: the Mars Desert Research Station (MDRS) operation can be more sustainable? It is evaluated how environmental, economic, social and personal sustainability issues are presented and is also reflected how MDRS activities would collaborate to Sustainable Development Goals (SDGs), proposing some guidelines to sustainability. It is also important to ask: the results related to Mars would be applied to build a self-sustainable habitats in Earth, mainly in areas affected by climate change as deserts and semiarid regions as can be seen Brazilian Northeast? Reviewing the previous research done at MDRS not was identified any research related to sustainability. Because of that, this research presents a possible great impact to MDRS. It is a challenger identifies the main dimensions that would be considered to evaluates a Mars research station in terms of sustainability and would be considered the great relevance of this research for the future of design of process of future Mars settlements.

The research Sustainability in MDRS: lessons to Mars research habitats and setlements was considered in the Mission Plan prepared today.

19h – CapComm